OpenStudy (anonymous):

At the beginning of an experiment, a scientist has 204 grams of radioactive goo. After 135 minutes her sample decayed to 12.75 grams. What is the half life? I just need help setting up the equation my book failed to do an example on this.

OpenStudy (dumbcow):

half life is time it takes to decay to 102 grams or half the initial amount. Pt = 204*e^-kt first solve for k 12.75 = 204*e^-135k